OpenStudy (anonymous):

One year ago Jesse was three times his sister's age. Next year he will be only twice her age. How old will Pat be in five more years?

OpenStudy (anonymous):

A) 8 B) 12 C) 11 D) 13 E) 15

OpenStudy (anonymous):

J=3S J+1=2(S+1) J+1=2S+2 sub 3S+1=2S+2

OpenStudy (anonymous):

Let j and p be Jesse and Pat's age today, respectively. Solve the following simultaneous equations for j and p: (j - 1) = 3 (p - 1), (j + 1) = 2 (p + 1) Then add 5 to the value for p. p=8, choice A)